11. In printing an article of 48,000 words, a printer decides to use two sizes of type. Using the
larger type, a printed page contains 1,800 words. Using a smaller type, a page contains
2,400 words. The article is allotted 21 full pages in a magazine. How many pages must be
in smaller type? ___________

Found 2 solutions by richwmiller, pattymint:
Answer by richwmiller(17219)

l+s=21
1800l+2400s=48000
l=4 s=17

Answer by pattymint(1)

Let x = the pages in small print.
21-x = the pages in large print
2400x + 1800(21-x) = 48,000
2400x + 37,800 - 1800x = 48,000
600x = 10,200
x = 17pages in small print
21-17 = 4 pages in large print.
Check your work!
2400(17) + 1800(4) = 48,000
40,800 + 7200 = 48,000
48,000 = 48,000